OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2.

OpenStudy (anonymous):

\[\sqrt{2}*\sqrt[3]{2}\]

OpenStudy (anonymous):

Is it like that?

OpenStudy (anonymous):

Change each radical to rational expression. \[\sqrt{2}-->2^{\frac{ 1 }{ 2 }}\] \[\sqrt[3]{2}-->2^\frac{ 1 }{ 3 }\] Then use LCD in both fractions.

OpenStudy (anonymous):

i have no idea how to do it. I think its 2 2/3

OpenStudy (anonymous):

and yes it looks like that

OpenStudy (anonymous):

What is the least common multiple of 2 and 3?

OpenStudy (anonymous):

2: 2, 4, 6, 8, 10 3: 3, 6, 9, 12, 15

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Yes! So,